C=100+0.5yd,Tax=100,I=200,Export=120,G=150,Import=0.03Y. Calculate the equilibrium net income and net exports


1
Expert's answer
2021-03-17T18:32:47-0400

"Y=C+I+G+Export-Import"

"0.53Y=100+200+150+120"

"Y=1075.5"

the equilibrium net income


"I+T+Export-Import-C=-249.75"

net exports


"Export-Import=120-32=88"